OpenStudy (anonymous):

Solve. 36 – 3x = 9 A. –12 B. –9 C. 9 D. 15

OpenStudy (anonymous):

9

OpenStudy (texaschic101):

36 - 3x = 9 -3x = -36 + 9 -3x = - 27 x = -27/-3 x = 9 Night-Watcher is correct

OpenStudy (anonymous):

Thank you @Night-Watcher

OpenStudy (anonymous):

You to @texaschic101

OpenStudy (anonymous):

No problem.

OpenStudy (texaschic101):

thats what we are here for :)